OpenStudy (anonymous):

The radius of one sphere is twice as great as the radius of a second sphere. a. Find the ratio of their surface areas. b. Find the ratio of their volumes.

ganeshie8 (ganeshie8):

lets say, radius of one sphere : r radius of second sphere : 2r surface area = 4 . pi . r^2 ratio of surface area of one sphere to the surface area of second sphere : \[\frac{4 . \pi . r^{2}}{4 . \pi . (2r)^{2}}\]

ganeshie8 (ganeshie8):

can you simplify above ratio ?